# All-electrically controlled spintronics in altermagnetic heterostructures

### v2

@@ III.1 Scattering formalism and spin-polarized conductance
[…]
@@
with $\Omega=\sqrt{(J^{2}-t^{2})a^{2}k_{y}^{2}+t_{\sigma}(E-U_{g})}\text{,}$ and $t_{\sigma}=t+\sigma J\cos(2\theta_{J})$, which depends on the strength $J$, orientation $\theta_{J}$, and the gate $U_{g}$ applied in AM, as well as the incident angle $\theta_{k}$ and energy $E$ of the injecting electron. Using the boundary conditions associated with Eq. (III.1), we impose
@@
and $\alpha=\pm 1$. The resulting spin-resolved transmission probability is
which determines the spin-resolved conductance at zero temperature [82]
where $G_{0}=e^{2}Wk_{F}/(2\pi h)$ is the conductance unit, and $W$ is the sample width. The total conductance of the junction is given by the sum over spin channels,
$G=G_{\uparrow}+G_{\downarrow},$ (21)
where the labels $\uparrow$ and $\downarrow$ correspond to $\sigma=+1$ and $\sigma=-1$, respectively. In the linear region, the spin-resolved current driven by an applied bias $V_{\text{bias}}$ is $I_{\sigma}=G_{\sigma}V_{\text{bias}}$ and total current is $I=I_{\uparrow}+I_{\downarrow}$ [82]. To quantify the contribution of each spin channel, we define the spin polarization as
$P=\frac{G_{\uparrow}-G_{\downarrow}}{G_{\uparrow}+G_{\downarrow}},$ (22)
where $P>0$ ($P<0$) indicates a dominant contribution from spin-up (spin-down) electrons. The spin polarization $P$ serves as a measure of spin-filtering efficiency when a spin-degenerate electron beam is injected. In particular, $P=+1$ ($P=-1$) corresponds to a fully spin-polarized conductance and current carried entirely by spin-up (spin-down) states. The generation and electrical control of spin-polarized conductance and thus the current are essential goals in the development of spintronic devices.
Figure 3: (a) Schematic illustration of the spin-filter effect. (a-i) The spin-filter is composed of normal metals and a gated altermagnet by $U_{g}$. The junction is along the $x$-direction, keeping periodic boundary conditions along the $y$-direction. (a-ii): The normal-metal leads exhibit spin-degenerate parabolic dispersions (black lines), while the weak (strong) AM features spin-dependent parabolic bands with the same (opposite) curvature direction, with blue and red lines denoting the spin-up and spin-down states, respectively. The dispersions are exhibited along $k_{x}$ direction with $k_{y}=0$. (a-iii) and (a-iv): Isotropic Fermi surface of normal-metal leads and anisotropic closed and open spin-resolved Fermi surfaces for weak and strong AMs, respectively. The Fermi surfaces are exhibited in $k_{x}$-$k_{y}$ plane. Blue (red) shaded regions indicate transverse mode ($k_{y}$) ranges forbidden for spin-up (spin-down) electrons due to the AM band structure. (b) Transmission probability, spin-resolved conductance and spin polarization for a weak AM. (b-i) and (b-iii): Energy and incident-angle ($\theta_{k}$) resolved transmission probabilities $T_{\uparrow}$ and $T_{\downarrow}$. (b-ii) and (b-iv): Transmission for normal incidence ($\theta_{k}=0$) at various tunneling barrier heights $V$. The horizontal dashed line denotes the resonant energy levels [Eq. (24)] due to the confinement condition [Eq. 23]. (b-v) and (b-vi): Energy-dependent spin-resolved conductance and resulting spin polarization. (c) Same as panel (b), but for a strong AM. Parameters: $U_{L}=0$, $U_{g}=1$, $V=0$ in (i, iii, v), and junction length $d=20a$. The $d_{x^{2}-y^{2}}$-wave AM with $\theta_{J}=0$ is considered.
III.2 Transport in weak altermagnetic junctions
Based on the scattering formalism above, in this section, we analyze transport through a spin filter based on a weak $d_{x^{2}-y^{2}}$ [Fig. 3(a-ii) and (a-iii)], characterized by a closed, spin-dependent Fermi surface [Fig. 2(a)]. Figure 3(b) presents the energy- and angle-resolved transmission probabilities, as defined in Eq. (19), along with the resulting spin-resolved conductance [Eq. (20)] and spin polarization [Eq. (22)].
The spin-dependent transmission probabilities $T_{\uparrow}$ and $T_{\downarrow}$ are shown in Figs. 3(b-i) and 3(b-iii), respectively, as functions of the incident energy $E$ and injection angle $\theta_{k}$. The corresponding normal-incidence transmission profiles, $T_{\uparrow}(\theta_{k}=0)$ and $T_{\downarrow}(\theta_{k}=0)$, are plotted in Figs. 3(b-ii) and 3(b-iv) for different values of the tunneling barrier height $V$. Resonant transmission ($T_{\sigma}=1$) is achieved when the junction length $d$ satisfies a standing wave condition [76] such that it equals an integer multiple of half the wavelength of the spin-$\sigma$ electron in the altermagnetic region, where $\lambda_{\sigma}=2\pi/k_{+}^{\sigma}$. This results in the resonance condition,
$k_{+}^{\sigma}d=n\pi,\quad(n=1,2,3,\cdots),$ (23)
which ensures constructive interference between the left-going and right-going propagating modes in AM in Fabry–Pérot–like structures [76, 77, 78, 79]. So that the phase difference accumulated over distance $d$ between the two modes is an integer multiple of $\pi$. With this condition, one can check that the denominator of Eq. (17) becomes minimal, resulting in maximal transmission probability. By substituting Eq. (15) into the resonance condition [Eq. (23)], the corresponding resonant energy levels related to the confined standing wave are
with $t_{\pm}=t\pm\sigma J$. The resonant energy levels depend on the spin index $\sigma$, the transverse incident angle $\theta_{k}$, and device parameters such as $U_{L}$, $U_{g}$, and $d$, but are independent of the tunneling barrier $V$. Thus, the transmission probability remains unity even in the presence of a finite tunneling barrier $V$ when the incident energy satisfies the resonance condition, $E=E_{n}^{\sigma}(\theta_{k})$, but becomes strongly suppressed in the off-resonant regime, $E\in(E_{n}^{\sigma},E_{n+1}^{\sigma})$ [See Fig. 3(b-i)-(b-iii) for $n=1$]. The energy spacing between neighboring resonant levels is
By summing over all possible incident directions, the spin-resolved conductance [Fig. 3(b-v)] exhibits the spin-dependent resonance condition described by Eq. (23). Specifically, the spin-up conductance is significantly suppressed by the tunneling barrier except for the resonant peaks, while the spin-down conductance remains measurable. These resonant features correspond to constructive interference conditions for the spin-dependent wavefunctions inside the altermagnetic region. As the tunneling barrier height $V$ increases, the off-resonant transmission is further suppressed, thereby enhancing the spin polarization $P$ [Fig. 3(b-vi)]. However, because of the continuous and overlapping band structure of the weak AM, both spin channels inevitably contribute to transport for a given energy, see Fig. 3(b-i) and (b-iii). Electrons incident at oblique angles ($\theta_{k}\neq 0$) retain finite tunneling probabilities. This leads to a residual spin-up contribution that limits the maximum achievable spin polarization. Consequently, even in the presence of a high tunneling barrier, full spin polarization ($P=1$ or $-1$) is hardly achieved. Instead, the spin polarization saturates at approximately $P\sim 80\%$ for large $V$, as demonstrated in Fig. 3(b-vi).
The partial polarization in weak AM discussed above poses two notable challenges for device applications: (i) The maximum spin polarization is not fully tunable and remains sensitive to the Fermi energy $E$. This limits the flexibility of weak AM-based spin filters compared to conventional ferromagnetic systems [12, 13], where full spin polarization can be realized and externally modulated through magnetic fields [83] or spin-transfer torques [84]. (ii) Enhancing spin filtering requires a high tunneling barrier, which inherently reduces the total conductance and the current. The inverse relationship between spin polarization and conductance imposes a fundamental constraint on device performance, especially in regimes where a sizable spin-polarized current is required for detection or functionality. While the challenge (i) can be mitigated by applying gate voltages to shift the resonance conditions as demonstrated in Sec. IV.1, the challenge (ii) is intrinsic to the weak AM band structure. However, as we show in the following, this limitation can be circumvented by employing strong AM materials. In that case, the emergence of spin-polarized transport arises from the intrinsic band structure of the strong AM itself, rather than relying on interfacial tunneling effects. This results in robust and fully spin-polarized transport, even in the absence of a tunneling barrier.
III.3 Transport properties in strong altermagnetic Junctions
The distinct dispersion of a strong AM [Fig. 2(b)] enables a fundamentally different electrically controlled transport behavior compared to its weak counterpart. In this case, we note that the resonant conditions in Eqs. (23)–(III.2) still apply. In particular, the spin polarization of the Fermi surface in a strong AM can be switched by tuning the energy, as illustrated in Fig. 3(a-iv), which is a behavior absent in weak AMs. As a result, spin-down (spin-up) electrons with $\theta_{k}^{-}<\theta_{c}^{-}$ ($\theta_{k}^{+}<\theta_{c}^{+}$) are blocked for $E>U_{g}$ ($E<U_{g}$), as shown in Fig. 3(c-ii) and Fig. 3(c-iii), respectively. This can also be understood from the possible Fermi surface contours at positive energy $E$, illustrated in the inset of Fig. 2(b). In this regime, spin-up (spin-down) states are absent when $a|k_{x}|<\sqrt{(E-U_{g})/(t+J)}$ ($a|k_{y}|<\sqrt{(E-U_{g})/(t+J)}$). Conversely, at negative energies, spin-up (spin-down) states vanish when $a|k_{y}|<\sqrt{(E-U_{g})/(t-J)}$ ($a|k_{x}|<\sqrt{(E-U_{g})/(t-J)}$), see Fig. 3(a-iv).
III.4 Effects of the altermagnetic orientation $\theta_{J}$
IV Gate-Controlled Spin Filtering and Spin Valve Effect in Altermagnet junctions
IV.1 Gate-controlled spin filter effects
In the weak AM regime, the conductance exhibits spin-dependent oscillations as the gate potential $U_{g}$ increases from zero [Fig. 4(a)], arising from quantum confinement effects [77, 78]. The conductance peaks appear when the resonance condition [Eq. (23)] is satisfied. Between the resonance peaks, one spin channel is strongly suppressed by the tunneling barrier, while the other remains finite, leading to asymmetric spin-resolved conductances. This asymmetry gives rise to finite spin polarization $P$ [Fig. 4(b)]. At low barrier strength (e.g., $V=0$), the conductance is only partially spin-polarized, whereas at higher barriers (e.g., $V=10$), a fully spin-polarized conductance is realized with a switchable polarization between $P=+1$ and $P=-1$ by tuning the gate voltage $U_{g}$.
IV.2 Electrically controlled spin valve based on strong AMs
@@
To explain the transport mechanism within this electrically modulated spin valve, we use Eq. (III.1) but with an electrostatic potential in AM given by
@@
By solving the scattering problem using the same procedure as in Sec. III.1, the spin-resolved transmission $T_{\sigma}$ in the spin valve configuration is obtained. Using Eqs. (20)–(22), the total conductance $G$ and the resulting spin polarization $P$ are evaluated. Their dependence on the gate voltages $U_{g1}$ and $U_{g2}$ is shown in Figs. 5(b) and 5(c), respectively.
Figure 5(b) demonstrates the switching between high and low conductance states by tuning the relative values of the gate voltages $U_{g1}$ and $U_{g2}$ in the spin valve based on a strong AMs [Fig. 5(a)]. This effect originates from the opposite spin-dependent band-opening directions in strong AMs [Fig. 5(a)]. As a result, transmission through each gated region is dominated by spin-up (spin-down) electrons for positive (negative) gate voltages, consistent with the spin-filtering behavior shown in Fig. 4(d). A high-conductance state arises when $U_{g1}U_{g2}>0$, indicating aligned spin polarizations in the two gated AM regions [Fig. 5(b)]. This configuration is analogous to the parallel alignment of magnetizations in conventional ferromagnetic spin valves [13, 14]. More specifically, for $U_{g1},U_{g2}>0$ ($<0$), transport is fully carried by spin-up (spin-down) carriers, yielding a spin polarization of $P=+1$ ($-1$), as shown in Fig. 5(c). In contrast, when $U_{g1}U_{g2}<0$, the opposite spin polarization in the two gated regions leads to a strong suppression of conductance due to spin-filter mismatch, resulting in an ill-defined $P$ through Eq. (22). This behavior is reminiscent of the antiparallel configuration in conventional spin valves, where spin-polarized electrons are blocked by the misaligned magnetic orientation of the second layer [13, 14].
V Conclusions
We have investigated spin-selective transport in heterostructures composed of altermagnetic materials and normal-metal leads, focusing on the contrasting behaviors between weak and strong altermagnets. By analyzing the interplay between altermagnetic exchange interaction and kinetic energy, we classified altermagnets into weak and strong regimes, each exhibiting distinct spin-resolved Fermi surface configurations. These differences manifest in qualitatively distinct transport signatures. While weak altermagnets support partial spin filtering via resonant tunneling enhanced by interface barriers, strong altermagnets intrinsically support fully spin-polarized conductance due to the absence of one spin species near the Fermi surface. Building upon these findings, we proposed an electrically tunable spin valve based on strong altermagnets. In this device, double-gate control allows transitions between parallel and antiparallel spin configurations without requiring magnetic fields or net magnetization, unlike conventional ferromagnetic spin valves. This functionality stems from the nonrelativistic spin-splitting inherent to strong AMs and highlights their potential as a platform for scalable, all-electrical spintronic devices. Our results demonstrate that altermagnets, particularly in the strong regime, offer a robust and versatile route to realizing spintronic functionalities through purely electric control. This work paves the way for future experimental efforts in designing gate-tunable spin filters and spin valves in candidate materials such as weak altermagnets RuO2 [40, 23, 25] and strong altermagnets Mn5Si3 [26, 42].
Acknowledgements.
@@
P.-H. Fu and Q. Lv contribute equally to this work. P.-H. Fu appreciates the support from W. Xu and the discussion from K. W. Lee and Z. Yuan. Y. Xu acknowledges financial support from the Scientific Research Starting Foundation of Ningbo University of Technology (Grant No. 2022KQ51). J. C. acknowledges financial support from the Carl Trygger’s Foundation (Grant No. 22: 2093), the Sweden-Japan Foundation (Grant No. BA24-0003), the Göran Gustafsson Foundation (Grant No. 2216), and the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121). J.-F. L. acknowledges financial support from the National Natural Science Foundation of China (Grant No. 12174077) and the Joint Fund with Guangzhou Municipality under Grant No. 202201020238. X.-L.Yu acknowledges financial support from the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515011852).

### v3

@@ Methods
and the notation $\left\{\cdots\right\}$ in the third line of Eq. (Methods) indicates the anticommutator, ensuring the Hermiticity of the Hamiltonian [86, 96]. Periodic boundary conditions are assumed along the $y$ direction, so that $k_{y}$ is conserved throughout the three regions of the junction.
@@
with $\Omega=\sqrt{(J^{2}-t^{2})a^{2}k_{y}^{2}+t_{\sigma}(E-U_{g})}\text{,}$ and $t_{\sigma}=t+\sigma J\cos(2\theta_{J})$, which depends on the strength $J$, orientation $\theta_{J}$, and the gate $U_{g}$ applied in AM, as well as the incident angle $\theta_{k}$ and energy $E$ of the injecting electron. Using the boundary conditions associated with Eq. (Methods), we impose
@@
and $\alpha=\pm 1$. The spin-resolved transmission probability [Eq. 8] can be obtained from Eq. (26).
@@
To explain the transport mechanism within this electrically modulated spin valve shown in Fig. 5(a), we use Eq. (Methods) but with an electrostatic potential in AM given by
@@
By solving the scattering problem using the same procedure, the spin-resolved transmission $T_{\sigma}$ in the spin valve configuration shown in Fig. 5(a) is obtained. Using Eqs. (9)–(11), the total conductance $G$ and the resulting spin polarization $P$ are evaluated. Their dependence on the gate voltages $U_{g1}$ and $U_{g2}$ is shown in Figs. 5(b) and 5(c), respectively.
Data Availability
The data supporting the findings of this study are available from the corresponding authors upon reasonable request.
Acknowledgements
@@
P.-H. Fu appreciates the support from W. Xu and the discussion from K. W. Lee and Z. Yuan. Q. Lv acknowledges financial support from Shenzhen University of Information Technology (Grant No. SZIIT2025KJ066). Y. Xu acknowledges financial support from the Scientific Research Starting Foundation of Ningbo University of Technology (Grant No. 2022KQ51) and the China Postdoctoral Science Foundation (Grant No. 2023M743783). J. C. acknowledges financial support from the Carl Trygger’s Foundation (Grant No. 22: 2093), the Sweden-Japan Foundation (Grant No. BA24-0003), the Göran Gustafsson Foundation (Grant No. 2216), and the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121). J.-F. L. acknowledges financial support from the National Natural Science Foundation of China (Grant No. 12174077) and the Joint Fund with Guangzhou Municipality under Grant No. 202201020238. X.-L.Yu acknowledges financial support from the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515011852) and the Shenzhen Natural Science Foundation (Grant No. JCYJ20250604174400001).
Author Contributions
P.-H. F. and Q. L. contributed equally to this work. P.-H. F. conceived the idea, developed the theoretical model, performed calculations, and wrote the manuscript. Q. L. also performed calculations and contributed to manuscript writing. X. Y. and J. C. provided valuable insights and contributed to the analysis and interpretation of the results. X.-L. Y. and J.-F. L. supervised the project. All authors discussed the results and contributed to the final version of the manuscript.
Competing Interests
The authors declare no competing interests.